OpenStudy (anonymous):

standard form for a line parallel to the line y=5x-12 and passing through the point (-2,-6)

OpenStudy (anonymous):

if you are looking for a line that is parallel, the slope will be the same. y = 5x + 12.....your slope is 5. now we will use point slope formula : y - y1 = m(x - x1) using slope(m) = 5 (-2,-6) x1 = -2 and y1 = -6 now lets sub our stuff in... y - (-6) = 5(x - (-2) y + 6 = 5(x + 2) y + 6 = 5x + 10 y = 5x + 10 - 6 y = 5x + 4 standard form is Ax + By = C y = 5x + 4 (subtract 5x from both sides) -5x + y = 4 (multiply by -1 to make x positive) 5x - y = -4 any questions ?

OpenStudy (anonymous):

No, thank you very much your awesome.